# Label-efficient Time Series Representation Learning: A Review

### v3

@@ Abstract
The scarcity of labeled data poses a significant challenge to deploying deep learning models for time series data in real-world applications. To address this, various strategies, e.g., transfer learning, self-supervised learning, and semi-supervised learning, have been developed to enhance the learning capabilities of these models using limited labeled time series data. In this survey, we introduce a novel taxonomy for the first time, categorizing existing approaches based on their reliance on external data sources to tackle the labeled data scarcity issue in time series. Furthermore, we present a review of the recent advances in each strategy, conclude the limitations of current methodologies, and suggest future research directions that promise further improvements in the field.
@@ I Introduction
  • We evaluate the challenges, emerging trends, and future directions in label-efficient time series analysis.
@@ II Related Surveys
Regarding the cross-domain related surveys, inspired by the success of transfer learning in computer vision, Ismail Fawaz et al. [11] explored the potential of transfer learning for time series classification and conducted exhaustive experiments to draw conclusions. Their findings underscore the presence of transferable low-level features in time series, and their transferability can be measured by the Dynamic Time Warping technique. Last, Shi et al. [5] offer a thorough discussion about unsupervised domain adaptation in the context of time series sensor data. They categorize these studies by industry, detailing application backgrounds, frequently used sensors, data discrepancy sources, and prominent research datasets.
@@ III-A Literature Search and Selection Strategy
Search Strategy: We utilized leading academic databases, including IEEE Xplore, PubMed, Google Scholar, and ACM Digital Library, focusing on publications in the last five years. Keywords, e.g., “time series analysis”, “deep learning”, and “semi-supervised learning” guided our search, aiming to include a wide range of studies pertinent to label efficiency in time series data.
@@ III-B Representation Learning Problem Formulation
In this section, we present a mathematical framework that encapsulates the essence of time series label-efficient learning, covering both in-domain and cross-domain learning scenarios. This framework will set the stage for a deeper exploration of the specific strategies you mentioned.
Consider a time series dataset $\mathcal{D}$ composed of pairs $(X,Y)$, where $X=\{x_{1},x_{2},...,x_{N}\}$ represents the time series data points and $Y=\{y_{1},y_{2},...,y_{N}\}$ denotes their corresponding labels. In label-efficient learning scenarios, the goal is to maximize the predictive performance of a learning model $f:X\rightarrow Y$ while minimizing the dependence on labeled instances. Given that a time series instance $x_{i}\in X$ is a sequence of observations $x_{i}=\{o_{1},o_{2},...,o_{T}\}$, where $T$ is the length of the series and $o_{t}$ represents the observation at time $t$. For supervised scenarios, each $x_{i}$ has a corresponding label $y_{i}\in Y$. In label-efficient scenarios, $Y$ may be partially observed or observed with noise.
In-Domain Learning: The dataset $\mathcal{D}$ can vary in composition, encompassing labeled $\mathcal{D}_{L}=\{(x_{i},y_{i})\}_{i=1}^{L}$, unlabeled $\mathcal{D}_{U}=\{x_{j}\}_{j=L+1}^{N}$, or a combination of both subsets. The learning objective adapts to this variability, aiming to utilize $\mathcal{D}_{L}$, $\mathcal{D}_{U}$, or their integration to effectively train a model $f$ that can accurately predict $Y$ for new instances, accommodating the dataset’s specific labeling scenario.
@@ III-C1 Convolutional Neural Networks (CNNs)
CNNs, including 1D-CNN [14], 1D-ResNet [15], and temporal convolutional network (TCN) [12], are prevalent in time series classification. Other designs were proposed, introducing different designs, e.g., InceptionTime [16] and Omni-Scale CNN blocks [17]. These networks excel at identifying patterns and trends with fewer parameters, reducing overfitting risks.
@@ III-C3 Attention Mechanism
The attention mechanism was proposed to enhance LSTM by focusing on specific input parts and assigning scores to hidden states [21]. Later, it became vital in applications [22, 23, 24]. Recently, the Transformer model became emerging as a prominent architecture for time series [25, 26]. While attention improves long sequence handling and model interpretability, its computational cost and overfitting risk due to more parameters are notable drawbacks.
@@ IV-A Data Augmentation
Figure 2: Increasing the number of samples through augmentations. For each input sample, we apply three (or more) augmentations to generate new samples
@@ IV-B Self-supervised Learning
Figure 3: Pipeline of the self-supervised learning process. The pretraining phase receives the input time series, sends it to the encoder to extract the features, and performs the self-supervised learning task. Next, the pretrained encoder is fine-tuned with the available few labeled data.
@@ IV-B2 Contrastive-based Methods
Contrastive-based self-supervised learning methods aim to form positive and negative pairs, where the contrastive loss maximizes the similarity among positive pairs while minimizing the similarity among negative pairs. To illustrate, given an unlabeled dataset $\mathcal{D}_{U}=\{x_{j}\}_{j=1}^{N}$ of time series instances, the objective of contrastive self-supervised learning is to learn a representation $f$ that embeds similar instances, i.e., positive pairs closer together and dissimilar instances, i.e., negative pairs, further apart in the embedding space. The positive pairs are data instances $(x_{i},x_{i}^{+})$ derived from the same time series instance or through augmentations that preserve the semantic content (see Section IV-A). The negative pairs are $(x_{i},x_{j}^{-})$, where $x_{i}$ and $x_{j}^{-}$ do not share the same semantic content, often chosen from different time series instances within the dataset. The objective is to learn a function $f$ that minimizes a contrastive loss, such as the InfoNCE loss, defined for a batch of instances as follows:
$\mathcal{L}_{\text{contrastive}}=-\sum_{i=1}^{n}\log\frac{\exp(sim(f(x_{i}),f(x_{i}^{+}))/\tau)}{\sum_{x_{j}^{-}\in N}\exp(sim(f(x_{i}),f(x_{j}^{-}))/\tau)},$
where $sim(u,v)$ is a similarity measure between two embeddings (e.g., the dot product or cosine similarity), $\tau$ is a temperature parameter that scales the similarity measure. Notably, the sum in the denominator runs over a set of negative pairs for each positive pair, and $n$ is the number of positive pairs in a batch.
@@
ExpCLR [41] leverages expert knowledge instead of relying on data transformations for contrastive learning. Recognizing the availability of expert features in fields like industry and medicine, this method incorporates such features to ensure the development of useful time-series representations that fulfill two proposed essential properties often overlooked by current methods. Mixup Contrastive Learning (MCL) proposes to replace the traditional contrastive loss with a mixup-based contrastive loss [42]. Specifically, MCL generates new mixed samples in each mini-batch using the mixup operation, and the proposed loss predicts the mixup ratio among each pair of samples. Also, Chang et al. [43] propose another loss that searches for positive and negative pairs for each anchor sample in both the input and embedding spaces. The contrastive loss is then applied to the selected triplet to learn a causal temporal CNN.
@@ IV-C Semi-supervised Learning
Semi-supervised learning is a type of learning that involves training a model using both labeled and unlabeled data. Combining a small labeled dataset $\mathcal{D}_{L}$ with a larger unlabeled dataset $\mathcal{D}_{U}$, the goal is to train a model $f$ that leverages the unlabeled data $\mathcal{D}_{U}$ alongside $\mathcal{D}_{L}$ to enhance learning. The learning objective often combines a supervised loss $\mathcal{L}_{\text{sup}}(f(x_{i}),y_{i})$ on $\mathcal{D}_{L}$ with an unsupervised loss $\mathcal{L}_{\text{unsup}}(f(x_{j}))$ on $\mathcal{D}_{U}$, where $\mathcal{L}_{\text{unsup}}$ encourages the model to learn meaningful structures from $\mathcal{D}_{U}$.
Semi-supervised learning is particularly useful in situations where it is difficult or expensive to obtain a large amount of labeled data, and only a few labeled samples are available with a large amount of unlabeled data. One of the challenges in semi-supervised learning is how to achieve the best balance of learning from both labeled and unlabeled samples. Fig. 4 provides a visual illustration of the semi-supervised training process. In specific, we feed the model with both the labeled and unlabeled samples, apply the traditional cross-entropy classification loss to the labeled data, and train the model on the unlabeled data with an unsupervised learning technique. In this survey, we categorize semi-supervised learning approaches into self-training-based methods and generic regularization-based methods.
@@ IV-C1 Self-training Methods
Figure 4: Illustration of semi-supervised learning process. Usually, the classification cross-entropy loss is applied to the labeled portion of data, while an unsupervised learning technique is applied to learn useful and supportive representations from the unlabeled data.

### v4

@@ Abstract
Label-efficient time series representation learning, which aims to learn effective representations with limited labeled data, is crucial for deploying deep learning models in real-world applications. To address the scarcity of labeled time series data, various strategies, e.g., transfer learning, self-supervised learning, and semi-supervised learning, have been developed. In this survey, we introduce a novel taxonomy for the first time, categorizing existing approaches as in-domain or cross-domain, based on their reliance on external data sources or not. Furthermore, we present a review of the recent advances in each strategy, conclude the limitations of current methodologies, and suggest future research directions that promise further improvements in the field.
@@ I Introduction
  • We explore the challenges, emerging trends, and future directions in label-efficient time series analysis.
@@ II Related Surveys
Regarding the cross-domain related surveys, inspired by the success of transfer learning in computer vision, Fawaz et al. [11] explored the potential of transfer learning for time series classification and conducted exhaustive experiments to draw conclusions. Their findings underscore the presence of transferable low-level features in time series, and their transferability can be measured by the Dynamic Time Warping technique. Last, Shi et al. [5] offer a thorough discussion about unsupervised domain adaptation in the context of time series sensor data. They categorize these studies by industry, detailing application backgrounds, frequently used sensors, data discrepancy sources, and prominent research datasets.
@@ III-A Literature Search and Selection Strategy
Search Strategy: We utilized major academic databases, including IEEE Xplore, PubMed, Google Scholar, and ACM Digital Library, focusing on publications from the past five years. Our search was guided by carefully chosen keywords including “time series analysis”, “deep learning”, “self-supervised learning”, “semi-supervised learning”, “unsupervised domain adaptation”, “transfer learning”, “semi-supervised domain adaptation”, “weakly supervised learning”, “active learning”, and “prototypical networks”, each appended with “time series” or some related applications, e.g., “EEG”, “Human Activity Recognition”, and “fault diagnosis”, to focus specifically on time series contexts.
@@ III-B Representation Learning Problem Formulation
In this section, we present a mathematical framework that encapsulates the essence of time series label-efficient learning, covering both in-domain and cross-domain learning scenarios.
Consider a time series dataset $\mathcal{D}$ composed of pairs $(\mathbf{x},\mathbf{y})$, where $\mathbf{x}=\{x_{1},x_{2},...,x_{N}\}$ represents the collection of time series data points and $\mathbf{y}=\{y_{1},y_{2},...,y_{N}\}$ denotes their corresponding labels. In label-efficient learning scenarios, the goal is to maximize the predictive performance of a learning model $f:\mathcal{X}\rightarrow\mathcal{Y}$ while minimizing the dependence on labeled instances. Here, $\mathcal{X}$ and $\mathcal{Y}$ represent the input and output spaces of all possible time series instances, respectively. Each time series instance $x_{i}\in\mathbf{x}$ is a sequence of observations $x_{i}=\{o_{1},o_{2},...,o_{T}\}$, where $T$ is the length of the series and $o_{t}$ represents the observation at time $t$. For supervised scenarios, each $x_{i}$ has a corresponding label $y_{i}\in\mathbf{y}$. In label-efficient scenarios, $\mathbf{y}$ may be partially observed or observed with noise.
In-Domain Learning: The dataset $\mathcal{D}$ can vary in composition, encompassing labeled samples only $\mathcal{D}_{L}=\{(x_{i},y_{i})\}_{i=1}^{L}$ if $L=N$, unlabeled samples only $\mathcal{D}_{U}=\{x_{j}\}_{j=1}^{E}$ if $E=N$, or a combination of both subsets if $L+E=N$. The learning objective adapts to this variability, aiming to utilize $\mathcal{D}_{L}$, $\mathcal{D}_{U}$, or their integration to effectively train a model $f$ that can accurately predict $\mathbf{y}$ for new instances, accommodating the dataset’s specific labeling scenario.
@@ III-C1 Convolutional Neural Networks (CNNs)
CNNs, including 1D-CNN [14], 1D-ResNet [15], and temporal convolutional network (TCN) [12], are prevalent in time series classification. Other designs have been proposed, introducing different designs, e.g., InceptionTime [16] and Omni-Scale CNN blocks [17]. These networks excel at identifying patterns and trends with fewer parameters, reducing overfitting risks.
@@ III-C3 Attention Mechanism
The attention mechanism was proposed to enhance LSTM by focusing on specific input parts and assigning scores to hidden states [21]. Later, it was adopted by several time series applications [22, 23, 24]. Recently, the Transformer model became emerging as a prominent architecture for time series [25, 26]. While attention improves long sequence handling and model interpretability, its computational cost and overfitting risk due to more parameters are notable drawbacks.
For the following sections, the block “Encoder” in diagrams is generic and can be adopted as one of the above networks, allowing for flexibility depending on the specific requirements and constraints of the application.
@@ IV-A Data Augmentation
Figure 2: Data augmentation is utilized to increase the number of training samples. For each input sample, several augmentations are applied to generate new samples.
@@ IV-B Self-supervised Learning
Figure 3: The pipeline of the self-supervised learning process. In the pretraining phase, the encoder is trained with the self-supervised loss without labels. Next, in the fine-tuning phase, the pretrained encoder is fine-tuned with the available few labeled data.
@@ IV-B2 Contrastive-based Methods
Contrastive-based self-supervised learning methods aim to form positive and negative pairs, where the contrastive loss maximizes the similarity among positive pairs while minimizing the similarity among negative pairs. To illustrate, given an unlabeled dataset $\mathcal{D}_{U}=\{x_{j}\}_{j=1}^{N}$ of time series instances, the objective of contrastive self-supervised learning is to learn a representation $f$ that embeds similar instances, i.e., positive pairs closer together and dissimilar instances, i.e., negative pairs, further apart in the embedding space. The positive pairs are data instances $(x_{i},x_{i}^{+})$ derived from the same time series instance or through augmentations that preserve the semantic content (see Section IV-A). The negative pairs are $(x_{i},x_{j}^{-})$, where $x_{i}$ and $x_{j}^{-}$ do not share the same semantic content, often chosen from different time series instances within the dataset. The objective is to learn a function $f$ that minimizes a contrastive loss, such as the InfoNCE loss [13], defined for a batch of instances as follows:
$\mathcal{L}_{\text{contrastive}}=-\sum_{i=1}^{n}\log\frac{\exp(sim(f(x_{i}),f(x_{i}^{+}))/\tau)}{\sum_{x_{j}^{-}\in N}\exp(sim(f(x_{i}),f(x_{j}^{-}))/\tau)},$ (1)
where $sim(u,v)$ is a similarity measure between two embeddings (e.g., the dot product or cosine similarity), $\tau$ is a temperature parameter that scales the similarity measure. Notably, the sum in the denominator runs over a set of negative pairs for each positive pair, and $n$ is the number of positive pairs in a batch. In addition, the hyperparameter $\tau$ plays a crucial role in controlling the scale of the distribution of distances between embeddings. A smaller value of $\tau$ sharpens the distribution increasing the sensitivity to small differences in distances, potentially focusing the learning on hard negatives. Conversely, a larger $\tau$ value smooths out the distribution, which can mitigate the impact of outliers but may reduce discriminative power by treating dissimilar pairs more leniently.
@@
ExpCLR [41] leverages expert knowledge instead of relying on data transformations for contrastive learning. Recognizing the availability of expert features in fields like industry and medicine, this method incorporates such features to ensure the development of useful time-series representations that fulfill two proposed essential properties often overlooked by current methods, i.e., the proximity of representations for samples with similar expert features and the separation of representations for samples with markedly different expert features. These properties are essential in defining a new loss function tailored to handle continuous expert features. Mixup Contrastive Learning (MCL) proposes to replace the traditional contrastive loss with a mixup-based contrastive loss [42]. Specifically, MCL generates new mixed samples in each mini-batch using the mixup operation, and the proposed loss predicts the mixup ratio among each pair of samples. Also, Chang et al. [43] propose another loss that searches for positive and negative pairs for each anchor sample in both the input and embedding spaces. The contrastive loss is then applied to the selected triplet to learn a causal temporal CNN.
@@
Figure 4: Illustration of the semi-supervised learning process. We train the model with the labeled portion of the data via the cross-entropy loss $\mathcal{L}_{\text{cls}}$, and the unlabeled portion via an unsupervised learning technique $\mathcal{L}_{\text{unsup}}$, i.e., $\mathcal{L}_{\text{SemiSL}}=\mathcal{L}_{\text{cls}}+\mathcal{L}_{\text{unsup}}$.
@@ IV-C Semi-supervised Learning
Semi-supervised learning is a type of learning that involves training a model using both labeled and unlabeled data. It becomes beneficial when it is difficult or expensive to obtain a large amount of labeled data, and only a few labeled samples are available with a large amount of unlabeled data. Specifically, given a small labeled dataset $\mathcal{D}_{L}$ with a larger unlabeled dataset $\mathcal{D}_{U}$, the goal is to train a model $f$ that leverages the unlabeled data $\mathcal{D}_{U}$ alongside $\mathcal{D}_{L}$ to enhance learning. The learning objective often combines a supervised classification loss $\mathcal{L}_{\text{cls}}(f(x_{i}),y_{i})$ on $\mathcal{D}_{L}$ with an unsupervised loss $\mathcal{L}_{\text{unsup}}(f(x_{j}))$ on $\mathcal{D}_{U}$, where $\mathcal{L}_{\text{unsup}}$ encourages the model to learn meaningful structures from $\mathcal{D}_{U}$. As illustrated in Fig. 4, the overall objective becomes $\mathcal{L}_{\text{SemiSL}}=\mathcal{L}_{\text{cls}}+\mathcal{L}_{\text{unsup}}$, where both losses are added together during training. However, one challenge in semi-supervised learning still achieving the best balance of learning from both labeled and unlabeled samples.
In this survey, we categorize semi-supervised learning approaches into self-training-based methods and generic regularization-based methods.